Find two rational numbers between 1 and 999/1000
I know that 1 is the same as 1000/100, so I do not know how there can be a rational number between 1000/100 and 999/1000

Answer by Alan3354(69443)

Find two rational numbers between 1 and 999/1000
I know that 1 is the same as 1000/100, so I do not know how there can be a rational number between 1000/100 and 999/1000
-------------
1 = 1000/1000, not 1000/100.
------------
1 = 100000/100000
999/1000 = 9990/100000
--> 9991/100000, 9992/100000, etc